# Quantification of Interdependent Emotion Dynamics in Online Interactions

### v3

@@ Discussion
First, anger and disgust strongly reinforce one another. Their reciprocal triggering, along with high self-excitation, aligns with the role of these moral emotions in driving moral outrage. Previous work has shown that co-occurrence of anger and disgust is predictive of moral outrage [88]; our findings suggest this co-occurrence may be further amplified by mutual excitation. This mechanism could help explain the frequent escalation of outrage in online discourse and its potential contribution to polarization and social fragmentation [89, 90]. We note, however, that part of this effect may be due to the frequent co-labeling of anger and disgust in the same message (see Appendix A).
@@ Conclusions
Appendix A Summary Statistics of YouTube Live Chat Data
We show some additional summary statistics of our dataset. Once we filtered the data by removing inactive videos, we compiled 92,412 live chat messages from 397 videos, amassing a total of 780 hours of material and encompassing 27 distinct topics. The distribution of topics is shown in Figure A.1. The distribution of emotional events is shown in Figure A.2. We also show the distribution of the median time between each event arrival in Figure A.3. To exemplify our emotion labeling approach, we showcase representative live chat messages and predicted emotions in Table A.1. Since we label emotions non-exclusively, we show the co-occurrence patterns across 6 basic emotions in Figure A.4. In general, negative emotions have a higher tendency to co-occur with one another. In particular, anger and disgust have a relatively high co-occurrence.
To examine whether expressing different emotions requires systematically different levels of effort, we analyze the distribution of message lengths across emotion categories. As shown in Figure A.5, the variation in text length within each category is substantially larger than the differences between categories. This suggests that observed asymmetries in emotional dynamics between positive and negative emotions are unlikely to be driven by differences in text length or typing effort.
Figure A.1: Number of videos per keyword in our final data sample across 27 keywords.
Figure A.2: (left) Distribution of the number of events per video across 397 videos. (right) Same as left but broken down per emotion.
Figure A.3: (left) Distribution of the median time interval between each event in units of minutes for each video across 397 videos. (right) Same as left but broken down per emotion.
Figure A.4: Co-occurrence of emotions in our data. We normalize the number of co-occurrences for each emotion pair by the sum of occurrences of the two emotions in the live chat.
Figure A.5: Distribution of YouTube Live Chat message lengths across emotion categories. The whiskers indicate the 5th to 95th percentiles, with both the median and mean marked within each box. The variation in text length within each emotion category is substantially larger than the differences in average text length across categories, suggesting that message length is not systematically biased by emotion type.
Table A.1: Sample live chat messages and detected emotions
@@ Appendix A Summary Statistics of YouTube Live Chat Data
Appendix B Baseline Intensity Calibration in Hawkes Process
@@
Appendix C Log-Likelihood Derivation
@@
We derive the log-likelihood function for a fairly general class of the Hawkes process as follows. The expected number of events for emotion $e$ within time interval $\Delta t$ is approximately $\lambda^{e}(t)\Delta t$ for small $\Delta t$’s. In the limit of an infinitesimally small time interval $\Delta t\to\mathrm{d}t$, the term $\lambda^{e}(t)\mathrm{d}t$ then represents the probability of observing exactly one event, since the probability of observing more than one event is an infinitesimal of higher order (technically proportional to $dt^{2}$). Similarly, the probability of not observing any event within $\mathrm{d}t$ reads $1-\lambda^{e}(t)\mathrm{d}t$, which, in the limit of $dt$ infinitesimal, becomes equivalent to $\exp\left(-\lambda^{e}(t)\mathrm{d}t\right)$. The probability of not observing any event over a finite time $\Delta t$ is thus $\exp\left(-\int\limits_{t}^{t+\Delta t}\leavevmode\nobreak\ \mathrm{d}\tau\leavevmode\nobreak\ {\lambda^{e}(\tau)}\right)$ by the law of independent probabilities. Taken together, the likelihood function capturing event intensity at $t_{i}$ over $[0,T]$ reads
$\prod_{i=1}^{N^{e}}\lambda^{e}(t_{i})e^{-\int\limits_{t_{i}}^{t_{i+1}}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\lambda^{e}(s)}}.$ (2)
Taking the product over the exponential terms, Equation (2) simplifies to $\prod_{i=1}^{N^{e}}\lambda^{e}(t_{i})e^{-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\lambda^{e}(s)}}$. The Hawkes process likelihood function is thus
$L^{e}(\theta^{e,f})=\prod_{i=1}^{N^{e}}\lambda^{e}(t_{i})e^{-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\lambda^{e}(s)}}.$ (3)
@@
$\log(L^{e}(\theta^{e,f}))$ $=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\lambda^{e}(s)}$ (4a)
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}(s)}-\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\phi^{e,f}(s)}.$ (4b)
@@
$\log L^{e}(\alpha^{e,f},\gamma^{e})$ =(4b)${=}\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}(s)}-\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\alpha^{e,f}\frac{1}{\gamma^{e}}e^{-\frac{1}{\gamma^{e}}(s-t_{j}^{f})}}$ (5a)
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}(s)}+\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}{\alpha^{e,f}e^{-\frac{1}{\gamma^{e}}(s-t_{j}^{f})}}\Big{|}_{t_{j}^{f}}^{T}$ (5b)
$=\sum_{i=1}^{N^{e}}\log\left(\mu^{e}(t_{i})+\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\alpha^{e,f}\frac{1}{\gamma^{e}}e^{-\frac{1}{\gamma^{e}}(t_{i}-t_{j}^{f})}\right)-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}(s)}+$
@@
The following sections involve variations of the baseline intensity function, whereas the endogenous cross-excitation component is unchanged. We thus define $\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\alpha^{e,f}\left(e^{-\frac{1}{\gamma^{e}}(T-t_{j}^{f})}-1\right)$ from Equation (5c) as $endo$ for the following derivation.
In our model, we allow for a coefficient term $\nu^{e,f}$ which scales the effect of emotion $f$ displayed in the video on emotion $e$ in the chat. Concretely, we model the cross-excitation effects of the baseline function for process $f$ on process $e$, namely $\mu^{e}(t)=\mu^{e}_{0}+\mu^{e}_{1}(t)\equiv\mu^{e}_{0}+\sum_{f\in\mathcal{E}}\leavevmode\nobreak\ \nu^{e,f}\leavevmode\nobreak\ S^{f}(t)$. We thus denote a multivariate Hawkes process for a given process as
@@
$\log(L^{e}(\mu^{e}_{0},\nu^{e,f},\theta^{e,f}))$ $=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\lambda^{e}(s)}$ (7a)
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}_{0}}-\sum_{f\in\mathcal{E}}\nu^{e,f}\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {S^{f}(s)}-\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\phi^{e,f}(s)}.$ (7b)
@@
$\log L^{e}(\mu^{e}_{0},\nu^{e,f},\alpha^{e,f},\gamma^{e})$ =(7b)${=}\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}_{0}}-\sum_{f\in\mathcal{E}}\nu^{e,f}\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {S^{f}(s)}+endo$ (8a)
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}_{0}}-\sum_{f\in\mathcal{E}}\nu^{e,f}\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {S^{f}(s)}+endo$ (8b)
@@
$\quad\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}_{0}}-\sum_{f\in\mathcal{E}}\nu^{e,f}\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {S^{f}(s)}+endo$ (8c)
We compute $\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {S^{f}(s)}$ numerically as $M^{f}_{1}$. Without loss of generality, we simplify the log-likelihood function to
@@
Appendix D Testing Log-Likelihood Fits with Synthetic Data
We estimate parameter values that maximize the log-likelihood function (C) via Quasi-Newton optimization. Without loss of generality, we normalize the log-likelihood function by the number of events $N^{e}$, in order to compare the log-likelihood values across different sample sizes. We use the Python ticks library 333 https://x-datainitiative.github.io/tick/index.html to simulate Hawkes processes with parameters that we refer to as the ground-truth. Subsequently, we compare the estimated parameters with their ground-truth values to evaluate performance. Since the parameters for each emotion are estimated separately, without loss of generality, we only show the synthetic test results for one emotion (process). The other processes have comparable performances. To assess the robustness of the estimation, we bootstrap the data and repeat the estimation 100 times.
Figure D.1: We plot the shape of the log-likelihood function in Equation (C) varying the target parameter while holding the other parameters fixed at their ground-truth values. The ground-truth values of the target parameter are indicated by the red dots. The log-likelihood values are normalized by the number of events $N^{e}$ to be comparable across varying numbers of data points. The inset plots show the distribution of the fitted parameters from 100 bootstrapped samples. The red dashed lines indicate the ground-truth values for the parameter of interest. We show results for the first process in a 6-dimensional multivariate Hawkes process. The other processes have comparable performances.
The inset plots of Figure D.1 shows the distributions of estimated $\mu^{1}_{0}$, $\gamma^{1}$, $\alpha^{1,f}$, and $\nu^{1,f}$ values for 100 bootstrap samples with $N^{e}$ around 10,000. The dashed red lines mark the ground-truth values for $\mu^{1}_{0}$, $\gamma^{1}$, $\alpha^{1,f}$, and $\nu^{1,f}$ respectively. The estimated parameters are reasonably robust and are centered on the ground-truth values. In addition, we plot the log-likelihood functions with respect to changes in the parameter value. We observe that the log-likelihood functions are well-behaved and maximized by the ground-truth parameter values, with the exception of $\gamma$. The value of the logarithmic likelihood function is insensitive to changes in $\gamma^{1}$ over a certain range, displaying sloppiness [105].
Appendix E Robustness Checks

### v4

@@ Discussion
First, anger and disgust strongly reinforce one another. Their reciprocal triggering, along with high self-excitation, aligns with the role of these moral emotions in driving moral outrage. Previous work has shown that co-occurrence of anger and disgust is predictive of moral outrage [85]; our findings suggest this co-occurrence may be further amplified by mutual excitation. This mechanism could help explain the frequent escalation of outrage in online discourse and its potential contribution to polarization and social fragmentation [86, 87]. We note, however, that part of this effect may be due to the frequent co-labeling of anger and disgust in the same message (see Supplementary Note 1).
@@
Another promising direction for future work is to reframe the definition of what is considered endogenous and exogenous while applying exactly the same modeling framework. In our current study, we treat the entire audience collectively as the endogenous system, with video content providing the primary exogenous input. Alternatively, one could define each individual viewer as the endogenous component and treat all other viewers and the video as exogenous influences. This reframing would not require any changes to the methodology itself: the same Hawkes process framework and estimation procedures can be directly applied, simply interpreting the resulting parameters at the individual rather than group level. Such an approach would allow for a richer understanding of how much an individual’s emotional expressions are shaped by their own prior states versus external social and media influences.
@@ Conclusions
Acknowledgements
The authors received no specific funding for this work.
Author Contributions
S.C.L., D.S., and Y.L. designed the research; Y.L. aggregated the data and wrote the code; S.C.L. D.S., and Y.L. analyzed the results; and S.C.L., D.S., and Y.L. wrote and reviewed the manuscript.
Competing Interest
The authors declare no competing interests.
Supplementary Information for
Social feedback amplifies emotional language in online video live chats
Appendix 1 Summary Statistics of YouTube Live Chat Data
We show some additional summary statistics of our dataset. Once we filtered the data by removing inactive videos, we compiled 92,412 live chat messages from 397 videos, amassing a total of 780 hours of material and encompassing 27 distinct topics. The distribution of topics is shown in Supplementary Figure 1. The distribution of emotional events is shown in Supplementary Figure 2. We also show the distribution of the median time between each event arrival in Supplementary Figure 3. To exemplify our emotion labeling approach, we showcase representative live chat messages and predicted emotions in Supplementary Table 1.
Since we label emotions non-exclusively, we show the co-occurrence patterns across 6 basic emotions in Supplementary Figure 4. In general, negative emotions have a higher tendency to co-occur with one another. In particular, anger and disgust have a relatively high co-occurrence.
To examine whether expressing different emotions requires systematically different levels of effort, we analyze the distribution of message lengths across emotion categories. As shown in Supplementary Figure 5, the variation in text length within each category is substantially larger than the differences between categories. This suggests that observed asymmetries in emotional dynamics between positive and negative emotions are unlikely to be driven by differences in text length or typing effort.
Supplementary Figure 1: Number of videos per keyword in our final data sample across 27 keywords.
Supplementary Figure 2: (left) Distribution of the number of events per video across 397 videos. (right) Same as left but broken down per emotion.
Supplementary Figure 3: (left) Distribution of the median time interval between each event in units of minutes for each video across 397 videos. (right) Same as left but broken down per emotion.
Supplementary Figure 4: Co-occurrence of emotions in our data. We normalize the number of co-occurrences for each emotion pair by the sum of occurrences of the two emotions in the live chat.
Supplementary Figure 5: Distribution of YouTube Live Chat message lengths across emotion categories. The whiskers indicate the 5th to 95th percentiles, with both the median and mean marked within each box. The variation in text length within each emotion category is substantially larger than the differences in average text length across categories, suggesting that message length is not systematically biased by emotion type.
Supplementary Table 1: Sample live chat messages and detected emotions
@@ Appendix 1 Summary Statistics of YouTube Live Chat Data
Appendix 2 Baseline Intensity Calibration in Hawkes Process
@@
Appendix 3 Log-Likelihood Derivation
@@
We derive the log-likelihood function for a fairly general class of the Hawkes process as follows. The expected number of events for emotion $e$ within time interval $\Delta t$ is approximately $\lambda^{e}(t)\Delta t$ for small $\Delta t$’s. In the limit of an infinitesimally small time interval $\Delta t\to\mathrm{d}t$, the term $\lambda^{e}(t)\mathrm{d}t$ then represents the probability of observing exactly one event, since the probability of observing more than one event is an infinitesimal of higher order (technically proportional to $dt^{2}$). Similarly, the probability of not observing any event within $\mathrm{d}t$ reads $1-\lambda^{e}(t)\mathrm{d}t$, which, in the limit of $dt$ infinitesimal, becomes equivalent to $\exp\left(-\lambda^{e}(t)\mathrm{d}t\right)$. The probability of not observing any event over a finite time $\Delta t$ is thus $\exp\left(-\int\limits_{t}^{t+\Delta t}\penalty 10000\ \mathrm{d}\tau\penalty 10000\ {\lambda^{e}(\tau)}\right)$ by the law of independent probabilities. Taken together, the likelihood function capturing event intensity at $t_{i}$ over $[0,T]$ reads
$\prod_{i=1}^{N^{e}}\lambda^{e}(t_{i})e^{-\int\limits_{t_{i}}^{t_{i+1}}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\lambda^{e}(s)}}.$ (7)
Taking the product over the exponential terms, Equation (7) simplifies to $\prod_{i=1}^{N^{e}}\lambda^{e}(t_{i})e^{-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\lambda^{e}(s)}}$. The Hawkes process likelihood function is thus
$L^{e}(\theta^{e,f})=\prod_{i=1}^{N^{e}}\lambda^{e}(t_{i})e^{-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\lambda^{e}(s)}}.$ (8)
@@
$\log(L^{e}(\theta^{e,f}))$ $=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\lambda^{e}(s)}$ (9a)
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\mu^{e}(s)}-\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\phi^{e,f}(s)}.$ (9b)
@@
[…]
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\mu^{e}(s)}+\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}{\alpha^{e,f}e^{-\frac{1}{\gamma^{e}}(s-t_{j}^{f})}}\Big|_{t_{j}^{f}}^{T}$ (10b)
$=\sum_{i=1}^{N^{e}}\log\left(\mu^{e}(t_{i})+\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\alpha^{e,f}\frac{1}{\gamma^{e}}e^{-\frac{1}{\gamma^{e}}(t_{i}-t_{j}^{f})}\right)-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\mu^{e}(s)}+$
@@
The following sections involve variations of the baseline intensity function, whereas the endogenous cross-excitation component is unchanged. We thus define $\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\alpha^{e,f}\left(e^{-\frac{1}{\gamma^{e}}(T-t_{j}^{f})}-1\right)$ from Equation (10c) as endoendo for the following derivation.
In our model, we allow for a coefficient term $\nu^{e,f}$ which scales the effect of emotion $f$ displayed in the video on emotion $e$ in the chat. Concretely, we model the cross-excitation effects of the baseline function for process $f$ on process $e$, namely $\mu^{e}(t)=\mu^{e}_{0}+\mu^{e}_{1}(t)\equiv\mu^{e}_{0}+\sum_{f\in\mathcal{E}}\penalty 10000\ \nu^{e,f}\penalty 10000\ S^{f}(t)$. We thus denote a multivariate Hawkes process for a given process as
@@
$\log(L^{e}(\mu^{e}_{0},\nu^{e,f},\theta^{e,f}))$ $=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\lambda^{e}(s)}$ (12a)
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\mu^{e}_{0}}-\sum_{f\in\mathcal{E}}\nu^{e,f}\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {S^{f}(s)}-\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\phi^{e,f}(s)}.$ (12b)
@@
[…]
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}))-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\mu^{e}_{0}}-\sum_{f\in\mathcal{E}}\nu^{e,f}\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {S^{f}(s)}+endo$ (13b)
@@
$\quad\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\mu^{e}_{0}}-\sum_{f\in\mathcal{E}}\nu^{e,f}\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {S^{f}(s)}+endo$ (13c)
We compute $\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {S^{f}(s)}$ numerically as $M^{f}_{1}$. Without loss of generality, we simplify the log-likelihood function to
@@
Appendix 4 Testing Log-Likelihood Fits with Synthetic Data
We estimate parameter values that maximize the log-likelihood function (3) via Quasi-Newton optimization. Without loss of generality, we normalize the log-likelihood function by the number of events $N^{e}$, in order to compare the log-likelihood values across different sample sizes. We use the Python ticks library (https://x-datainitiative.github.io/tick/index.html) to simulate Hawkes processes with parameters that we refer to as the ground-truth. Subsequently, we compare the estimated parameters with their ground-truth values to evaluate performance. Since the parameters for each emotion are estimated separately, without loss of generality, we only show the synthetic test results for one emotion (process). The other processes have comparable performances. To assess the robustness of the estimation, we bootstrap the data and repeat the estimation 100 times.
Supplementary Figure 6: We plot the shape of the log-likelihood function in Equation (3) varying the target parameter while holding the other parameters fixed at their ground-truth values. The ground-truth values of the target parameter are indicated by the red dots. The log-likelihood values are normalized by the number of events $N^{e}$ to be comparable across varying numbers of data points. The inset plots show the distribution of the fitted parameters from 100 bootstrapped samples. The red dashed lines indicate the ground-truth values for the parameter of interest. We show results for the first process in a 6-dimensional multivariate Hawkes process. The other processes have comparable performances.
The inset plots of Supplementary Figure 6 shows the distributions of estimated $\mu^{1}_{0}$, $\gamma^{1}$, $\alpha^{1,f}$, and $\nu^{1,f}$ values for 100 bootstrap samples with $N^{e}$ around 10,000. The dashed red lines mark the ground-truth values for $\mu^{1}_{0}$, $\gamma^{1}$, $\alpha^{1,f}$, and $\nu^{1,f}$ respectively. The estimated parameters are reasonably robust and are centered on the ground-truth values. In addition, we plot the log-likelihood functions with respect to changes in the parameter value. We observe that the log-likelihood functions are well-behaved and maximized by the ground-truth parameter values, with the exception of $\gamma$. The value of the logarithmic likelihood function is insensitive to changes in $\gamma^{1}$ over a certain range, displaying sloppiness [105].
Appendix 5 Robustness Checks